OpenStudy (anonymous):

4x - 2y. if x= -1 and y= -3

OpenStudy (anonymous):

69

OpenStudy (anonymous):

2

OpenStudy (anonymous):

4x - 2y = 4(-1) -2(-3) =-4 +6 =6-4=2

OpenStudy (uri):

4(-1)-2(-3) -4+6 =2